Love Theatre Inc. owns and operates movie theaters throughout New Mexico and Utah. Love Theatre has declared the following annual dividends over a six-year period: 2011, $20,000; 2012, $60,000; 2013, $95,000; 2014, $110,000; 2015, $140,000; and 2016, $180,000. During the entire period ending December 31 of each year, the outstanding stock of the company was composed of 25,000 shares of cumulative, 2% preferred stock, $100 par, and 100,000 shares of common stock, $5 par.

Required:

1. Calculate the total dividends and the per-share dividends declared on each class of stockfor each of the six years. There were no dividends in arrears on January 1, 2011. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0".

Preferred Dividends Common Dividends

Year
Total
Dividends

Total

Per Share

Total

Per Share
2011 $   20,000 $ $ $ $
2012 60,000
2013 95,000
2014 110,000
2015 140,000
2016 180,000
$ $

2. Calculate the average annual dividend per share for each class of stock for the six-year period. If required, round your answers to two decimal places.

Average annual dividend for preferred $ per share
Average annual dividend for common $ per share

3. Assuming a market price per share of $105 for the preferred stock and $11 for the common stock, calculate the average annual percentage return on initial shareholders' investment, based on the average annual dividend per share for preferred stock and for common stock.

Round your answers to two decimal places.

Preferred stock %
Common stock %
Answer #1

Since the preferred shares are cumulative, unpaid dividend of preferred stockholders would be accumulated and paid before anything is paid to common stockholders. The preference dividend of earlier years will also be paid before current year dividend of common stockholders.

Dividend per annum of preferred stockholders = 25,000*100*2% = $50,000

Year

Total Dividend

Preferred Dividends

Common Dividends

Total

Per Share

Accumulated

Total

Per Share

2011

20,000

20,000

0.8

30,000

0

0

2012

60,000

60,000

2.4

20,000

0

0

2013

95,000

70,000

2.8

0

25,000

0.25

2014

110,000

50,000

2

0

60,000

0.6

2015

140,000

50,000

2

0

90,000

0.9

2016

180,000

50,000

2

0

130,000

1.3

2.Annual Average Dividend for preferred = (0.8+2.4+2.8+2+2+2)/6

= $2 per share

Annual average dividend for common = (0+0+0.25+0.6+0.9+1.3)/6

= $0.5083 per share

3.Annual average percentage return = Annual average dividend/market price per share

Preferred Stock = 2/105 = 1.90%

Common Stock = 0.5083/11 = 4.62%